OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (25, 30) and (125, 50), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.) My answer is 1.5 Am I right? :)

OpenStudy (amistre64):

2 steps you can follow, subtract one point from the other; then stack y over x (125,50) -(25,30) -------- 100, 20 ; slope = 20/100 which is definantly less than 1

OpenStudy (amistre64):

is that . spose to be a / ?

OpenStudy (anonymous):

I think so

OpenStudy (amistre64):

20/100 = 1/5 so it looks like you hit the "." instead of the "/" key :)

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

It was supposed to be 1/5

OpenStudy (amistre64):

then yes, that is correct

OpenStudy (anonymous):

Think you so much! :)